### v2

# Expressing and visualizing model uncertainty in Bayesian variable selection using Cartesian credible sets

Abstract
Modern regression applications can involve hundreds or thousands of variables which motivates the use of variable selection methods. Bayesian variable selection defines a posterior distribution on the possible subsets of the variables (which are usually termed models) to express uncertainty about which variables are strongly linked to the response. This can be used to provide Bayesian model averaged predictions or inference, and to understand the relative importance of different variables. However, there has been little work on meaningful representations of this uncertainty beyond first order summaries. We introduce Cartesian credible sets to address this gap. The elements of these sets are formed by concatenating sub-models defined on each block of a partition of the variables. Investigating these sub-models allow us to understand whether the models in the Cartesian credible set always/never/sometimes include a particular variable or group of variables and provide a useful summary of model uncertainty. We introduce methods to find these sets that emphasize ease of understanding. The potential of the method is illustrated on regression problems with both small and large numbers of variables.
Keywords: Spike-and-slab prior; Posterior inclusion probability; median model; regression
1 Introduction
Regression methods allow us to relate the variation in a response to observed variables. If the number of variables are large then variable selection methods are needed. Their use is generally motivated by the need to avoid overfitting, which can reduce predictive performance, or to understand which subsets of the variables provide parsimonious explanations of the relationship with the response. Many methods have been proposed. Classical approaches include subset selection, stepwise selection methods or penalized maximum likelihood, including Lasso, elastic net, SCAD, or MCP (see Hastie et al., 2015, for a review). Bayesian methods place a prior distribution on the regression coefficients leading to a posterior distribution that can express the relative ability of different subsets of the variables to explain the variation in the response. This allows us to understand the relative importance of the variables and the relationship between the variables (for example, that two covariates carry similar information). The approach has been applied widely including linear models (Mitchell and Beauchamp, 1988; George and McCulloch, 1997), generalized linear models (Raftery, 1996) and nonparametric regression (Savitsky et al., 2011; Liu et al., 2021; Linero and Du, 2023).
There are two main classes of Bayesian approach. Firstly, global-local shrinkage priors encourage coefficients which explain little variation to be shrunk very close to zero (see e.g. Bhadra et al., 2019, for a review). Post-processing methods can be used to perform variable selection often by finding a sparse summary of the posterior distribution (Hahn and Carvalho, 2015; Piironen et al., 2020). Alternatively, spike-and-slab priors place positive probability on some regression coefficients being exactly zero (Mitchell and Beauchamp, 1988; George and McCulloch, 1997). This leads to a posterior distribution on all possible models defined by subsets of the variables. This approach has good theoretical properties (Castillo et al., 2015) and predictive performance (Porwal and Raftery, 2022). Recent work has provided efficient algorithms for use in high-dimensional problems involving tens of thousands of regressors, including Markov chain Monte Carlo (MCMC) methods such as Importance Tempering (Zanella and Roberts, 2019), ASI (Griffin et al., 2021), LIT (Zhou et al., 2022) and PARNI (Liang et al., 2022), or stochastic search methods such as SVEN (Li et al., 2023). We will follow the spike-and-slab approach using MCMC but the methods could be extended to approximations of the posterior from a spike-and-slab prior found using different computational methods such as stochastic search.
Bayesian variable selection leads to a posterior distribution on all possible models but inference or prediction usually averages over this posterior distribution to provide summaries or Bayesian model averaged predictions or estimates. The relative importance of the different variables is often assessed by the posterior probability that a variable is included in the model, which is usually termed the marginal posterior inclusion probability (PIP). The posterior distribution can also be summarized by a single model. The Maximum a posterior (MAP) model is the posterior modal model (Rockova and George, 2014) and the median model (Barbieri and Berger, 2004) includes all variables with a marginal PIP above 0.5. Methods that average over the posterior distribution are feasible even in large spaces using MCMC methods but accurate estimation of posterior model probabilities is usually infeasible if there are many variables.
Although, summaries can provide better estimates than a single model (see e.g. Barbieri et al., 2021; Porwal and Raftery, 2022), they do not allow us to effectively understand model uncertainty in general. Credible sets are a promising method for providing more information about model uncertainty in Bayesian variable selection. These guard against overinterpretation of the results of Bayesian variables from considering a single model. Initial work has been in the area of Genome Wide Association Studies (GWAS), where Single Nucleotide Polymorphisms (SNPs) are regressed on a trait of interest (such as disease status). Maller et al. (2012) proposed a simple method of finding the credible set for a single regression effect. Wang et al. (2020) extend this approach to multiple regression effects using the “sum of single effects” (SuSiE) approach. This assumes that there are $L$ non-zero regression effects and associates a distribution $\pi_{i}$ over the SNPs for the $i$-th regression effect, which summarises the uncertainty about the position of the effect. Wang et al. (2020) provide a method for finding a variational approximation to the posterior distribution of the $\pi_{i}$’s. The method is particularly effective in GWAS where SNPs are serially correlated due to linkage disequilibrium and it is assumed there is a fixed number of causal SNPs (non-zero regression coefficients). However, this approach has some drawbacks that can make it less suitable for general regression problems. Firstly, it assumes only negative correlation between the inclusion of variables and so is not suitable for forms of multi-collinearity which lead to positive correlation. Secondly, the value of $L$ must be chosen. Thirdly, credible sets are created for each of the $L$ regression effects rather than the model (which may include different numbers of variables).
In this paper, we take a post-processing approach, which has been successfully used in Bayesian regression (Hahn and Carvalho, 2015; Piironen et al., 2020), with a spike-and-slab prior and work directly on the posterior distribution of all possible models. We build credible sets on the whole model by approximating the posterior distribution as a product of blocks of variables and minimizing a criterion that balances the fidelity of the approximation with the complexity of credible sets for each block. This approach allows us to visualize the posterior distribution through the model uncertainty in each block and involves only quantities calculated as averages over models (so making it feasible using MCMC output).
The paper is organized as follows. Section 2 briefly reviews Bayesian variable selection. Section 3 introduces a method for constructing credible model sets with a factorized posterior. Section 4 discusses approximating the posterior distribution with a distribution factorized according to a partition of the variables, and choosing the partition to balance the fidelity of the approximation with the complexity of the credible set using an “ease-of-understanding” criterion. Section 5 describes visualization of the Cartesian credible sets. Section 6 applies the methods to a simulated example. Section 7 shows how the method can be used with a GWAS of Systemic Lupus Erythematosus using a logistic regression model, and a linear regression model with two-way interactions. Section 8 discusses. Appendices include proofs, additional details of the Bayesian methods used in the examples, and further results. Code to implement the method on MCMC samples is available from https://jimegriffin.github.io/website/.
2 Bayesian variable selection
Variable selection methods find a subset of $p$ possible variables which explain the variation in a response through a regression model. We refer to each possible combination of included variables as a model, which is represented by the $p$-dimensional binary inclusion vector $\gamma$ where the $i$-th variable is only included in the model if $\gamma_{i}=1$. Let $\Gamma=\{0,1\}^{p}$ be the space of all possible models. Bayesian approaches assume that a prior is placed on $\Gamma$ as well as any parameters of the regression model. Tadesse and Vannucci (2021) provide a book length review and discuss recent research directions. A main focus of interest is the posterior distribution of the models. In this paper, we do not make any assumptions about the form of regression model or priors used to construct this posterior and we will write it generically as $p(\gamma\mid\mbox{Data})$ for a model $\gamma\in\Gamma$. This is a discrete distribution over $2^{p}$ possible models and so is hard to visualize or even write down unless $p$ is very small.
The large size of $\Gamma$ also motivates the use of computational methods unless $p$ is small. There are two main computational approaches. Firstly, stochastic search methods where $p(\gamma\mid\mbox{Data})$ can be evaluated (for example, linear models) or approximated (for, generalized linear models) for a subset $\Gamma^{\star}$ of $\Gamma$. Several algorithms have been proposed to find a $\Gamma^{\star}$ which provides a good approximation to the posterior distribution (Hans et al., 2007; Li et al., 2023). The second approach uses MCMC to generate a sample $\gamma^{(1)},\dots,\gamma^{(N)}$ from the posterior distribution $p(\gamma|\mbox{Data})$. Unlike the first approach, this leads to unbiased estimates of posterior quantities but it can be computationally expensive to generate a representative sample (Zanella and Roberts, 2019; Griffin et al., 2021; Zhou et al., 2022; Liang et al., 2022).
As discussed in the introduction, there are several methods used to summarize the posterior distribution including: marginal PIPs, the Maximum a posterior (MAP) model and the median model (Barbieri and Berger, 2004; Barbieri et al., 2021). Barbieri and Berger (2004) show that the median model is optimal for prediction in orthogonal and nested design (when inclusion variables are independent a posteriori and the PIPs completely characterize the posterior distribution) and can be easily calculated from MCMC output. If there is correlation between variables then the inclusion variables will be a posteriori correlated and the PIPs do not summarize all information in the posterior. Barbieri et al. (2021) discuss properties of the median model under correlated designs. This is a particular problem if the inclusion variables have high correlation which corresponds to multi-collinearity in the design matrix of possible variables. To illustrate this effect, consider Example 5.2.2 in George and McCulloch (1997) where data is simulated from a linear regression model with $n=180$ observations and $p=15$ covariates. The only non-zero coefficients correspond to variables to 1, 3, 5, 7, 8, 11, 12, and 13. There is strong multicollinearity between the following blocks of variables: 1 and 2; 3 and 4; 5 and 6; 7, 8, 9, and 10; 11, 12 13, 14, and 15. Appendix B provides full details of the data generation and prior settings used in the analysis.
Correlation matrix of variables Correlation matrix of $\gamma$
Refer to caption Refer to caption
Figure 1: George and McCulloch (1997) example: the correlation matrix of the variables and the inclusion variables $\gamma$.
Figure 1 shows the correlation matrix of the variables and the posterior correlation matrix of the inclusion variables. The posterior correlation matrix of the inclusion variables shows blocks of correlated variables but with each block uncorrelated. The blocks unsurprisingly mirror the blocks of variables with strong multicollinearity. PIPs and the modal or median model cannot express this correlation information which show strong negative and positive associations. Negative associations are particularly interesting since these show that variables are interchangeable. For example, variables 1 and 2 carry very similar information (similarly for 3 and 4, and 5 and 6).
3 Cartesian credible sets with factorized posteriors
3.1 Credible sets
Credible sets extend credible intervals, which are commonly used to summarize posterior distributions, to posteriors on discrete spaces. Let $\mathcal{C}$ be the set of all possible subsets of $\Gamma$ then $S\subset\mathcal{C}$ is a $100\,\lambda\%$ credible set if $p(S\mid\mbox{Data})\geq\lambda$ (Wang et al., 2020). The $100\,\lambda\%$ highest posterior density credible interval is the shortest possible $100\,\lambda\%$ credible interval and, similarly, we can define the $100\,\lambda\%$ highest posterior probability (HPP) credible set, which is the smallest $100\,\lambda\%$ credible set. The $100\,\lambda\%$ HPP credible set can be found by ranking the models ranked list of the models in decreasing probability,
$p\left(\left.\gamma^{(1)}\right|\mbox{Data}\right)\geq p\left(\left.\gamma^{(2)}\right|\mbox{Data}\right)\geq p\left(\left.\gamma^{(3)}\right|\mbox{Data}\right)\geq\dots\geq p\left(\left.\gamma^{(2^{p})}\right|\mbox{Data}\right)$
and finding the smallest $K$ such that $\sum_{k=1}^{K}p\left(\left.\gamma^{(k)}\right|\mbox{Data}\right)\geq\lambda$ then $\left\{\gamma^{(1)},\gamma^{(2)},\dots,\gamma^{(K)}\right\}$ is the HPP credible set. This approach is possible if $p$ is relatively small (roughly less than 30) and all posterior probabilities can be calculated or approximated analytically. However, in many cases, $p$ will be larger and can be substantially larger (for example, genomic data sets can have tens of thousands of potential variables). In these cases, it’s often hard to accurately estimate model probabilities unless MCMC sample sizes are very large, as the level of model uncertainty makes it difficult to visit all plausible models, (i.e. models with posterior probability greater than $\epsilon$ for $10<\epsilon<$ 1 where, typically, $\epsilon$ is small) using stochastic search or MCMC. Additionally, searching the space of credible sets $\mathcal{C}$ is daunting as it contains $12^{2^{p}}-$ 1 possible sets and the size of the HPP credible set can make it hard to write down or consider.
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 Prob
1 0 0 1 1 0 0 0 1 1 0 0 0 1 1 0.0849
1 0 1 0 1 0 0 0 1 1 0 0 0 1 1 0.0817
1 0 0 1 1 0 0 0 1 1 1 1 1 0 0 0.0424
1 0 0 1 1 0 0 0 1 1 0 0 1 1 1 0.0396
0 1 1 0 1 0 0 0 1 1 0 0 0 1 1 0.0345
0 1 0 1 1 0 0 0 1 1 0 0 0 1 1 0.0338
1 0 1 0 1 0 0 0 1 1 0 0 1 1 1 0.0279
0 1 0 1 1 0 0 0 1 1 1 1 1 0 0 0.0264
0 1 0 1 1 0 0 0 1 1 0 0 1 1 1 0.0248
1 0 1 0 1 0 0 0 1 1 1 1 1 0 0 0.0218
1 0 1 0 0 1 0 0 1 1 0 0 0 1 1 0.0218
1 0 0 1 0 1 0 0 1 1 0 0 0 1 1 0.0202
0 1 1 0 1 0 0 0 1 1 0 0 1 1 1 0.0183
1 0 1 1 1 0 0 0 1 1 0 0 0 1 1 0.0176
0 1 1 0 1 0 0 0 1 1 1 1 1 0 0 0.0142
Table 1: George and McCulloch (1997) example: the models in the 50% HPP credible set shown with their posterior probability.. The vertical lines show the blocks of strongly multicollinear variables.
Before presenting a possible solution, consider the 50% HPP credible set for Example 5.2.2 of George and McCulloch (1997), which is shown in Table 1. There are 15 models but each model can be constructed by concatenating sub-models $\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$, $(0,0,1,1)$, and $\kappa_{4}$ to form $(\kappa_{1},\kappa_{2},0,0,1,1,\kappa_{4})$ where $\kappa_{1}\in\{(0,1),(1,0)\}$, $\kappa_{2}\in\{(0,1),(1,0)\}$, $\kappa_{3}\in\{(0,1),(1,0)\}$ and $\kappa_{4}\in\{(0,0,0,1,1),(1,1,1,0,0),(0,0,1,1,1)\}$. This suggests the following interpretation. Variables 9 and 10 should be included in the model and variables 7 and 8 excluded, only one of variables 1 and 2, one of variables 3 and 4, and one of variables 5 and 6 should be included in the model. Finally, one of the combinations: variables 11, 12 and 13, or variables 13, 14 and 15 or variables 14 and 15 should be included in the model. Expressing the models in the credible set as a concatenation of sub-models is easier to understand than enumerating all models in the 50% HPP credible set. There are 24 models that can be constructed in the form $(\kappa_{1},\kappa_{2},0,0,1,1,\kappa_{4})$ and this also defines a 50% credible set. This is larger than the 50% HPP credible set but much smaller than all possible models ($327682^{15}=$ 32768).
3.2 Finding Cartesian credible sets with factorized posteriors
The example in subsection 3.1 suggests that a simpler description of credible sets is possible if we restrict the space of possible credible sets from $\mathcal{C}$ to something more structured. A structure used throughout this paper is the Cartesian product of sets, which we will represent by $\otimes$, generated using a partition of the variables, which provides a formal way of concatenating elements of different sets.
Definition 1
Suppose that we have $p$ variables and a partition $P=\{P_{1},\dots,P_{L}\}$ of $\{1,2,\dots,p\}$ where $l_{i}$ is the size of the set $P_{i}$, $S=\otimes_{i=1}^{L}S_{i}\subset\Gamma$ where $S_{i}$ is a non-empty subset of $\{0,1\}^{l_{i}}$ for $i=1,\dots,L$ is a $100\lambda\%$ Cartesian credible set if $p(S\mid\mbox{Data})\geq\lambda$.
We will refer to each element of the partition as a block, $l_{i}$ as the size of the $i$-th block, and $S_{i}$ as a block credible set. It is also useful to define $\kappa_{i}\in\{0,1\}^{l_{i}}$ as a sub-model for the $i$-th block. This means that $S_{i}$ is a set of sub-models $\kappa_{i}$. The choice of a Cartesian product allows the credible set to be understood through the block credible sets in a similar way to our interpretation of the results from the George and McCulloch (1997) example above. Imposing this structure leads to $\prod_{i=1}^{L}(2^{2^{p_{i}}}-1)$ possible credible sets which is much smaller than all possible credible sets when the maximum of $p_{i}$ is much smaller than $p$. The Cartesian product is particularly useful when we assume that the posterior distribution can be factorized over the blocks using the following result.
Proposition 1
Suppose that $S=\otimes_{i=1}^{L}S_{i}$ and $p(\gamma\mid\mbox{Data})=\prod_{i=1}^{L}p(\kappa_{i}\mid\mbox{Data})$ for all $\kappa_{i}\in S_{i}$ for $i=1,\dots,L$ then
$p(S\mid\mbox{Data})=\prod_{i=1}^{L}p(S_{i}\mid\mbox{Data}).$
Example 1
Suppose that we have $p=3$ variables, the partition $P=\{1,\{2,3\}\}$ ($K=2$ blocks) and the posterior distribution in Table 2.
Block 1 Block 2
$\gamma_{1}=$ $(\gamma_{2},\gamma_{3})=$
0 1 (0, 0) (0, 1) (1, 0) (1, 1)
Posterior probability 0.9 0.1 0 0.5 0.5 0
Table 2: Posterior distribution for example 1.
If we choose $S_{1}=\{0,1\}$ and $S_{2}=\{(0,0),(0,1),(1,0),(1,1)\}$ then
$S=\{(0,0,0),(0,0,1),(0,1,0),(0,1,1),(1,0,0),(1,0,1),(1,1,0),(1,1,1)\}.$
The formula follows from noticing that $p(S)=1$, $p(S_{1})=1$ and $p(S_{2})=1$. Choosing $S_{1}=\{0\}$ and $S_{2}=\{(0,1),(1,0)\}$ leads to $S=\{(0,0,1),(0,1,0)\}$. The formula follows from noticing that $p(S)=0.9$, $p(S_{1})=0.9$, $p(S_{2})=1$.
Conveniently, we try to find small credible intervals but not necessarily the smallest. For example, the commonly used $100\lambda\%$ central credible interval is only the smallest $100\lambda\%$ credible interval if the posterior distribution is symmetric. Similarly, we would like to find a small $100\%\lambda$ Cartesian credible set. The smallest $100\,\lambda\%$ Cartesian credible set is the $S_{1},\dots,S_{L}$ which minimize $\sum_{i=1}^{L}\log\#S_{i}$ subject to $\sum_{i=1}^{L}\log P\left(S_{i}\mid\mbox{Data}\right)\geq\lambda$. By ordering the elements of $S_{i}$ to have decreasing posterior probability $p\left(\left.\kappa_{i}^{(1)}\right|\mbox{Data}\right)\leq\dots\leq p\left(\left.\kappa_{i}^{\left(2^{l_{i}}\right)}\right|\mbox{Data}\right)$ , this can be seen as a nonlinear integer knapsack optimisation problem which is known to be NP-hard in the case of arbitrary non-negative non-decreasing profit functions (Gurevysky et al., 2023). Gurevysky et al. (2023) derive approximate algorithms for this problem but these are not invariant to re-ordering of the blocks. This is not suitable here since we wish to treat all blocks in the same way. Therefore, we propose a different heuristic iterative algorithm. Suppose that we have current values $S_{1},\dots,S_{L}$, then removing the sub-model with the smallest marginal posterior probability from each block credible set leads to the smallest reduction in the probability of the Cartesian set, by a factor of $\frac{P\left(S_{i}\mid\mbox{Data}\right)-\min_{\kappa_{i}\in S_{i}}P(\kappa_{i}\mid\mbox{Data})}{\left(S_{i}\mid\mbox{Data}\right)}$. It also leads to the size of the Cartesian credible set reducing by a factor of #$\frac{\#S_{i}-1}{\#S_{i}}$. This suggests using a rule which balances removing a sub-model from the block which has the smallest effect on the probability of the Cartesian credible set with the largest effect on its size. This is achieved by choosing the block which leads to the smallest value of the criterion for the $i$-th block $\frac{\min_{\kappa_{i}\in S_{i}}P(\kappa_{i}\mid\mbox{Data})/\left(S_{i}\mid\mbox{Data}\right)}{1/\#S_{i}}$ and removing the sub-model with the smallest marginal posterior probability. This process is terminated when $\prod_{i=1}^{L}P\left(S_{i}\mid\mbox{Data}\right)<\lambda$ and the sets for which $\prod_{i=1}^{L}P\left(S_{i}\mid\mbox{Data}\right)\geq\lambda$ are used. It is convenient to define $\Gamma_{i}=\{0,1\}^{l_{i}}$ as the space of all sub-models in the $i$-th block and to sort the probabilities of the sub-models in ascending order. The algorithm is described in Algorithm 1.
$\pi_{i}\leftarrow 1$, $r_{i}=1$ and $q_{i}\leftarrow p\left(\left.\kappa^{\left(1\right)}_{i}\right|\mbox{Data}\right)$ for $i=1,\dots,L$.
while $\prod_{i=1}^{L}\pi_{i}\geq\lambda$ do
     $k^{\star}\leftarrow\arg\min(q_{1},\dots,q_{L})$.
     $\pi_{k^{\star}}\leftarrow\pi_{k^{\star}}-q_{k^{\star}}$
     $r_{k^{\star}}=r_{k^{\star}}+1$,
     $q_{k^{\star}}=\frac{p\left(\left.\kappa_{k^{\star}}^{\left(r_{k^{\star}}\right)}\right|\mbox{Data}\right)/\sum_{r=r_{k^{\star}}}^{2^{l_{i}}}p\left(\left.\kappa_{k^{\star}}^{\left(r_{k^{\star}}\right)}\right|\mbox{Data}\right)}{2^{l_{i}}-r_{k^{\star}}}$.
end while
Algorithm 1 Cartesian credible set finding algorithm
4 Cartesian credible sets with general posteriors
The method developed in Section 3 finds a $100\,\lambda\%$ Cartesian credible set if the posterior distribution can be factorized over the blocks in a partition. In many problems, the posterior distribution will not have a factor form and so we will choose a factorized approximation $q(\gamma|\theta)=\prod_{i=1}^{L}q(\kappa_{i}\mid\theta_{i})$ where $\theta=(\theta_{1},\dots,\theta_{L})$ are parameters. Papaspiliopoulos and Rossell (2017) consider how a block-diagonal design matrix leads to a factorised form of the posterior distribution and, in contrast to this paper, describe a block-diagonal approximation to the general design matrices. We first consider how to approximate the posterior distribution for a given partition and then how the partition can be chosen using an “ease-of-understanding” criterion which balances fidelity of the approximation with complexity of the credible set. In our approach, we exclude variables which have a marginal PIP below a threshold since these variables will not be included in the Cartesian credible set (we use a threshold of 0.04 in our examples).
The approximation $q$ for a particular partition is chosen by finding $\theta$ which minimize Kullback-Leibler divergence between the posterior and $q$,
KL $=\sum p(\gamma\mid\mbox{Data})\log p(\gamma)-\sum p(\gamma\mid\mbox{Data})\log q(\gamma)$
$=C-\sum_{\gamma\in\Gamma}p(\gamma\mid\mbox{Data})\log q(\gamma\mid\theta)$
$=C-\sum_{j=1}^{L}\sum_{\kappa_{j}\in\Gamma_{j}}p(\kappa_{j}|\mbox{Data})\log q(\kappa_{j}\mid\theta_{j}).$
If we can fully evaluate the posterior distribution then this can be calculated analytically. If there is an MCMC sample $\gamma^{(1)},\gamma^{(2)},\dots,\gamma^{(N)}\sim p(\gamma\mid\mbox{Data})$ then we can define $\kappa^{(i)}_{j}$ to be the values of $\gamma^{(i)}$ restricted to the block $P_{j}$. A Monte Carlo approximation to the KL divergence can be used
$-\frac{1}{N}\sum_{j=1}^{L}\sum_{i=1}^{N}\log q\left(\left.\kappa_{j}^{(i)}\right|\theta_{j}\right).$
In this paper, we define probabilities $q(\kappa_{i}=\phi)=\theta_{i}(\phi)$ for $\phi\in\Gamma_{i}$. The optimal choice under KL divergence is $\theta_{j}(\phi)=\frac{1}{N}\#\left\{\kappa_{j}^{(i)}=\phi\right\}$. Alternatively, we could use a parametric distribution such the quadratic exponential binary model (Cox and Weimurth, 1994) or the more general multivariate Bernoulli distribution (Dai et al., 2013).
We consider two algorithms with tuning parameters to find the partition which defines the factorized approximation. The first method uses a thresholded version of the correlation matrix and the second method applies an agglomerative algorithm to the KL divergence. In the thresholding method, we define a parameter $10<\eta<$ 1 and the matrix $\alpha(\eta)$ by $\alpha_{i,j}(\eta)=\mbox{I}(\|D_{i,j}|>\eta)$ for $i,j=1,\dots,p$ where $D$ is the posterior correlation matrix of $\gamma$. It is straightforward to show that $\alpha(\eta)$ is a well-defined adjacency matrix. The partition, as a function of $\eta$, is defined as the connected components of the graph defined by the adjacency matrix $\alpha(\eta)$. Clearly, the block sizes become larger as $\eta$ decreases. In practice, we find the partition for a grid of values of $\eta$, $01>\eta_{1}>\eta_{2}>\dots,\eta_{J}>$ 0 and find the corresponding partitions $P^{(1)},\dots,P^{(J)}$ and $100\lambda\%$ Cartesian credible set $S^{(1)},S^{(2)},\dots,S^{(J)}$ using Algorithm 1. In the agglomerative method, we define $\mbox{KL}(P)$ for a partition $P$ to be the Kullback-Leibler divergence between the posterior distribution and the approximation $q$ calculated using the optimal values of the parameters for $P$. A sequence of partitions can be calculated by merging two blocks of the partition at the previous iteration (in a similar way to agglomerative clustering). The two blocks to be merged are chosen to be ones that lead to the largest change in $\mbox{KL}(P)$. We define $\eta_{i}$ to be the change in $\mbox{KL}(P)$ at the $i$-th iteration of the algorithm. This greedy approach will lead to the largest improvement in the approximation at each stage and so defines a sequence of partitions $P_{1},\dots,P_{L}$ which are increasingly close approximation of the posterior distribution indexed by $\eta_{1},\dots,\eta_{L}$. The algorithm is summarized in Algorithm 2.
$P_{0}=\{\{1\},\{2\},\dots,\{p\}\}$.
for $k=1,\dots,p$ do
     Calculate $D_{i,j}=\mbox{KL}(P_{i,j})$ where $P_{i,j}$ is the partition joining the $i$-th and $j$-th element of $P$.
     Find $(i,j)$ which maximizes $D_{i,j}$ and let $a_{k}$ be this value.
     Set $P_{k}$ as $P_{k-1}$ with elements $i$ and $j$ joined and
     Set $\eta_{k}=KL(P_{k})-KL(P_{k-1})$.
     Set $S^{(k)}$ as the $100\lambda\%$ Cartesian credible set found using Algorithm 1
end for
Algorithm 2 Agglomerative algorithm
We could choose the tuning parameters using a criterion which measures the discrepancy between the approximation and the posterior (for example, using Kullback-Leibler divergence). However, this may lead to Cartesian credible sets with large block credible sets. Therefore, we use an alternative criterion which balances the fidelity of the approximation to the posterior distribution and the “ease-of-understanding” of the credible set. It is hard, or perhaps impossible, to provide a precise definition of “ease-of-understanding” of the credible set but we take the view that smaller block credible sets makes interpretation simpler and so the largest block credible set shouldn’t be “too large”. To achieve this, we use an “ease-of-understanding” penalized minimization criterion to the sequence of Cartesian credible sets define by $\eta_{1},\dots,\eta_{J}$. Our criterion is
$\log\#S^{(j)}-f(S^{(j)})$
where $f$ is a function of the credible set $S^{(j)}$ and penalizes partition leading to harder to understand credible sets. The first term measure the fidelity of the approximation since a partition with a smaller number of larger blocks will lead to better fidelity of the approximation and a smaller credible set. We choose to define the penalty on the partitions $P^{(j)}$ rather than the credible sets $S^{(j)}$. This is still challenging. We use a form inspired by the “rich-get-richer” property of the Dirichlet process (Ferguson, 1973) and penalizes partitions with large blocks. The penalty is the logarithm of the exchangeable product partition function of the Dirichlet process
$f(S^{(j)})=K^{(j)}\log M+\sum_{k=1}^{K^{(j)}}\log\Gamma\left(\#P_{i}^{(j)}\right).$
The parameter $M$ controls the level of penalization and we have found that $M=2$ works well in practice.
5 Visualizing the $100\lambda\%$ Cartesian credible set
The $100\lambda\%$ Cartesian credible set gives us a convenient way to visualize the posterior distribution of $\gamma$. Firstly, for each block in the partition, we can list the sub-models in the corresponding block credible interval. If a block only includes the null sub-model, ${\bf 0}_{i}=\underbrace{(0,\dots,0)}_{l_{i}\mbox{ times}}$, it’s excluded from the graph. To provide further information about the posterior distribution, we colour the sub-models according to their marginal PIP. We also colour the partition according the posterior probability that at least one variable in the block is included in the model which is $p(S_{i}-{\bf 0}_{i}\mid\mbox{Data})$. This colouring is useful for understanding whether the block provides evidence for the inclusion of one of these variables and allows us to address dilution effects (George, 2010), the splitting of posterior probability between models including one of a set of strongly correlated variables.
6 George and McCulloch (1997) Example
We find 50% Cartesian credible sets with simulated data generated using the set-up in Example 5.2.2 of George and McCulloch (1997) by applying the algorithms developed in Section 4.
Thresholding Method
Agglomerative Method
Figure 2: George and McCulloch (1997) Example with $M=2$. Black lines indicate the blocks of the partition.
The sets are shown in Figure 2. Both methods find partitions with similar blocks to those identified as having strong multicollinearity in Section 2. Both methods split $\{7,8,9,10\}$ into two blocks $\{7,8,10\}$ and $\{9\}$. This has no effect on the final result as variables 7 and 8 are excluded from the model and variables 9 and 10 are included. The thresholding method also splits $\{11,12,13,14,15\}$ into two blocks $\{11,12,14,15\}$ and $\{13\}$. This leads to a slightly larger set (32 models) compared with the Agglomerative Method (24 models). This compares to the smallest 50% credible set which has 15 models. The difference is small compared with the total number of models (32768).
The Cartesian credible sets allows us to identify the interchangeability of variables and the colouring allows us to understand the overall importance of a block of variables and the contribution of each sub-model to the overall importance. For example, $S_{1}=\{(0,1),(1,0)\}$ and the overall posterior probability of $S_{1}$ is close to 0.9, the sub-models in $S_{1}$ are $(0,1)$ and $(1,0)$ which have corresponding posterior probabilities 0.38 and 0.52. This can be interpreted as evidence that one of variable 1 or variable 2 should be included in the model but there is no evidence for preferring either variables (although, the posterior probability slightly favours variable 1). We can draw similar conclusions about the pairs variables 3 and 4, and variables 5 and 6. The interpretation of the variables 11 to 15 depends slightly on the method for finding the partition. Using the thresholding method, there is evidence to include only one of the pairs variables 11 and 12, or variables 14 and 15. There is no evidence to include or exclude variable 13. Using the agglomerative method, there is evidence to include at least one of the variables 11 to 15. Again, variables 11 and 12, and variables 14 and 15 act as pairs. The thresholding method gives a slightly simpler interpretation but also a larger set. Overall, these results give a more succinct expression of model uncertainty than the 50% HPP credible set in Table 1 and gives more information than the PIPs or median model.
7 Applications
7.1 GWAS example: Systemic Lupus Erythematosus
We consider a GWAS for Systemic Lupus Erythematosus using a case-control study. The presence or absence of the disease is the binary outcome in a logistic regression with the SNPs as explanatory variables. It is assumed that there are a number of causal SNPs which truly cause biological differences between the case and controls. A simple analysis would fit a logistic regression separately for each SNP and declare a SNP as associated with the outcome if it has a p-value smaller than a chosen (typically, small) threshold. However, due to correlation in the genome, we usually see many SNPs declared as associated for each causal SNP. We refer to this group of SNPs as the signal. The goal of the analysis is to determine the number of independent signals and which SNPs are most likely the causal SNPs at each signal.
Figure 3: GWAS example: Marginal PIPs for chromosome 3.
The data contains 4,036 cases and 6,959 controls, with the cases and controls having European ancestry. In this example we only look at SNPs on Chromosome 3 ($42,430$ SNPs/variables). We apply Bayesian variable selection to the data using a logistic regression model with all SNPs as potential covariates and fit the model using an adaptive MCMC algorithm (Griffin and Wan, 2021). The full specification of the Bayesian model is provided in Appendix C. Figure 3 shows the PIPs for each SNP. There are several clear signals from SNPs with PIPs equal to 1 and others with large PIPs. The 50% Cartesian credible sets are found by first retaining only those SNPs with a PIP above 0.04 (leaving 69 variables). The genomic distances of these 69 variables are given in Appendix C. The median model contains 13 variables which using the indices of the retained SNPs are 5, 10, 37, 45, 47, 48, 50, 51, 52, 53, 54, 58, 63.
Figure 4: GWAS example: Thresholding Method, $M=2$. Black lines indicate the blocks of the partition.
The 50% Cartesian credible sets found using the thresholding method and the agglomerative method are shown in Figures 4 and 5 respectively. The most striking aspect is that the blocks of variables are contiguous which reflects the serial correlation of the SNPs. The results using the thresholding method in Figure 4 show that 11 variables in the median model do not include the possibility of exclusion in the credible set (10, 45, 47, 48, 50, 51, 52, 53, 54, 58, 63). Two of these variables are included as a block: 58 with 55, 56 and 57, and 63 with 64, 65, 66 and 67. There are two blocks of variables which have no variables included in the median model but which have substantial support for inclusion as a block. These are 32, 33, 34, 35 and 36 (genomic distances: 103387440–103433922), which has a posterior probability of 0.61 of including one of the variables (the largest PIP for a variable in the block is 0.23), and 59, 60, 61, and 62 (genomic distance: 159767453–159791628), which has a posterior probability of 0.84 (the largest PIP for variables in the block is 0.28).
Figure 5: GWAS example: Agglomerative Method, $M=2$. Black lines indicate the blocks of the partition.
The results from the agglomerative method in Figure 5 are very similar to the results using the thresholding method but are more conservative with only the extra blocks $\{59,60,61,62\}$ and $\{63,64,65,66,67\}$ included. The method groups 57 with 58. The variables 32, 33, 34, 35 and 36 are not included as a block. The difference is marginal with the block $\{32,33,34,35,36\}$ from the thresholding method having fairly weak support to be included as a block.
7.2 Models with interactions: Concrete data
Schäfer and Chopin (2013) describe several data sets designed to have high-levels of multi-collinearity to test computational methods for sampling posterior distribution from Bayesian variable selection. We concentrate on their Concrete data example, which has 1030 observations and eight variables, which are listed in Appendix D. The reader is referred to their paper for full details of the regression problem. They construct a challenging variable selection problem by including transformation of the recorded covariates and two-way interactions. Table 3 shows the 13 variables created by transforming these original variables.
$X_{1}$ $\log X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $\log X_{4}$ $X_{5}$ $X_{6}$ $\log X_{6}$ $X_{7}$ $\log X_{7}$ $X_{8}$ $\log X_{8}$
1 2 3 4 5 6 7 8 9 10 11 12 13
Table 3: Concrete examples variables
Figure 6: Concrete data: Thresholding Method using prior I, $M=2$. Black lines indicate the blocks of the partition.
We consider how the choice of prior affects the model uncertainty through the lens of 50% Cartesian credible sets. We fit a linear regression which includes the 13 main effects and 78 two-way interactions, and consider two choices of prior. Firstly, we ignore the distinction between main effects and two-way interaction and treat this as a variable selection problem with 91 regressors. We use the hierarchical prior in Appendix B, which we term Prior I. An alternative prior can be constructed which distinguishes between main effects and interaction using the idea of strong heredity (that interactions can only be included in a model if both main effects are included). These types of prior were introduced in Chipman (1996). We use the prior in Appendix B but with the prior probability of 0.5 of including an interaction (conditional on including both main effects) and we give the prior of including a main effect, $\pi$, the prior $\pi\sim\mbox{Be}\left(1,\frac{p-3p_{0}+\sqrt{(p_{0}+p)^{2}+2pp_{0}(p-1)}}{2p_{0}}\right)$ where $p$ is the number of variables and $p_{0}$ is the prior mean number of included variables. The choice of prior for $\pi$ implies that the prior mean number of included variables is $pp_{0}$ as in Prior I. This is called Prior II. Both priors imply the same level of prior sparsity but differ in the structure of prior. With Prior I we used the MCMC set-up for the George & McCulloch example described in Appendix B and for Prior II we used an extension of the ASI algorithm (Griffin et al., 2021) to regression models with two-way interactions described in Appendix E
Figure 7: Concrete data: Thresholding Method using prior II, $M=2$. Black lines indicate the blocks of the partition.
The results of using the thresholding method with Priors I and II are shown in Figures 6 and 7 (the results using the Agglomerative Method are provided in Appendix D). The results with the two priors are strikingly different. Prior I leads to a much larger 50% Cartesian credible set which includes $3.98\times 10^{21}$ models whereas Prior II leads to a set with $4.95\times 10^{13}$. The set using Prior I is 80 000 000 times larger than the set using Prior II. This shows the ability of a well chosen prior to substantially reduce the amount of model uncertainty. Prior II also leads to much clearer results with an inclusion variable equal to 1 for 25 variables compared to 10 variables under Prior II.
8 Discussion
Cartesian credible sets are a useful way to summarize and visualize the output from Bayesian variable selection. The sets are constructed by finding a partition of the variable and an approximation to the posterior distribution where variables in different subsets are independent. Credible sets are defined as a Cartesian products of sets of sub-model and can be found using a simple heuristic algorithm. Two methods for finding the approximation are proposed: 1) thresholding the correlation matrix and 2) an agglomerative algorithm for minimizing the KL divergence between the posterior and the approximation. In both cases, there is a parameter controlling the level of approximation and this can be chosen using a criterion which trades-off the accuracy of the approximation and overall the size of the credible set. The method can be run quickly as post-processing of an MCMC sample from the posterior distribution on the models. The results can be visualized by plotting the list of sub-models for each subset and coloured according to the posterior probability of the sub-models and the posterior probability of including at least one variable in a subset. This gives a much more complete picture of the posterior distribution than the usually plotted PIPs and helps understanding of the interchangeability of variables which are highly collinear. This allows us to avoid dilution effects from posterior mass being distributed across such variables by considering blocks of variables.
The approach suggests several different directions for future research in Bayesian variable selection. The approximation of the posterior distribution from Bayesian variable selection using independent blocks of variables can be useful in computational methods. Some recently developed algorithms for high-dimensional variable selection use an independent approximation to the posterior distribution. In MCMC methods, the ASI algorithm (Griffin et al., 2021), and its development to ARNI and PARNI (Liang et al., 2022) use an independent approximation. Similarly, Ray and Szabó (2022) study variational inference for Bayesian variable selection where the variational distribution is independent across the variables. In both cases, the methods developed in this paper can act as starting point for defining iterative algorithms which learn the main features of the correlation structure in the posterior distribution. In this paper, we argue that a Cartesian credible set is a more useful summary of the posterior distribution than the smallest credible set. However, if the smallest credible set is interesting and the probability of individual models can be well-approximated by MCMC output, a Cartesian credible set could be used as a starting point. The smallest credible set can then be found by iteratively removing model with the smallest posterior probability.
Summarizing Bayesian variable selection is challenging as the posterior distribution is defined on a discrete space. Defining the credible set as a Cartesian product of sub-models on blocks of variables combines with an independent approximation of the posterior distribution to provide a simple summary method. There are many other Bayesian methods where the variable of interest is high-dimensional and discrete. For example, in Bayesian nonparametrics, clustering using mixture models and feature allocation or change-point models in time series. If an independent approximation is appropriate in these models, then the methods developed in this paper could be easily applied. In posterior distributions with more structure, other representations such as factor models could be used to define credible sets in these more complicated settings.
Acknowledgements
I would like to thank Xitong Liang and Dr Sam Livingstone from University College London and Professor Mike Smith from Melbourne Business School for helpful discussions. I would also like to thank Professor Tim Vyse and Dr David Morris from King’s College London for providing the data in the GWAS example and initial motivation to consider effectively expressing model uncertainty.
Appendix A Proofs
A.1 Proposition 1
$p(S|\mbox{Data})$ $=\sum_{\gamma\in S}p(\gamma|\mbox{Data})=\sum_{\gamma\in\otimes_{i=1}^{L}S_{i}}\prod_{i=1}^{L}p(\kappa_{i}|\mbox{Data})=\sum_{\kappa_{1}\in S_{1}}\dots\sum_{\kappa_{L}\in S_{L}}\prod_{i=1}^{L}p(\kappa_{i}|\mbox{Data})$
$=\prod_{i=1}^{L}\sum_{\kappa_{i}\in S_{i}}p(\kappa_{i}|\mbox{Data})=\prod_{i=1}^{L}p(S_{i}|\mbox{Data})$
Appendix B George and McCulloch (1997) example
This example has $n=180$ and $p=15$. Let $Z\sim\mbox{N}(0,I_{p})$ and $Z_{0}\sim\mbox{N}(0,1)$ and the variables are $X_{1}=Z_{1}+2Z_{0}$, $X_{3}=Z_{3}+2Z_{0}$, $X_{5}=Z_{5}+2Z_{0}$, $X_{8}=Z_{8}+2Z_{0}$, $X_{9}=Z_{9}+2Z_{0}$, $X_{10}=Z_{10}+2Z_{0}$, $X_{12}=Z_{12}+2Z_{0}$, $X_{13}=Z_{13}+2Z_{0}$, $X_{14}=Z_{14}+2Z_{0}$ $X_{15}=Z_{15}+2Z_{0}$, $X_{2}=X_{1}+0.15Z_{2}$, $X_{4}=X_{3}+0.15Z_{4}$, $X_{6}=X_{5}+0.15Z_{6}$, $X_{7}=X_{8}+X_{9}-X_{10}+0.15Z_{7}$, and $X_{11}=X_{14}+X_{15}-X_{12}-X_{13}+0.15Z_{11}$. The regression coefficients are $\beta=(1.5,0,1.5,0,1.5,0,1.5,-1.5,0,0,1.5,1.5,1.5,0,0)^{\prime}$ and so the non-zero regression coefficients are 1, 3, 5, 7, 8, 11, 12, 13.
We fit a linear regression model where $\alpha$ represents the intercept, $\beta_{\gamma}$ represents the regression coefficients for model $\gamma$, and $\sigma^{2}$ represents the observation variance. The prior distributions are
$p(\alpha,\sigma^{2})\propto\sigma^{-2},\quad\beta_{\gamma}\sim\mbox{N}(0,g\,I_{p+1}),\quad\gamma_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mbox{Bernoulli}(\pi),\ i=1,\dots,p,$
[…]
The posterior distribution was sampled using a simple Add-Delete-Swap sampler for $\gamma$ with Gibbs steps to update $\pi$ and $g$ and all other parameters integrated out. The samplers were run for 60 000 iterations with the first 10 000 iterations used as a burn-in, and the subsequent 50 000 iterations thinned every fifth value to leave a final sample of 10 000 values.
Appendix C GWAS data
We fit a logistic regression model where $\alpha$ represents the intercept, and $\beta_{\gamma}$ represents the regression coefficients for model $\gamma$. The prior distributions are
$\beta_{\gamma}\sim\mbox{N}(0,I_{p+1}),\quad\gamma_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mbox{Bernoulli}(\pi),\ i=1,\dots,p,\quad\pi\sim\mbox{Beta}\left(1,\frac{p-p_{0}}{p_{0}}\right),$
The posterior distribution was sampled using the ASI sampler for logistic regression from Griffin and Wan (2021). Three independent ASI chains were run with each chain containing 25 replicate chains. Each ASI was run for 15 000 iterations with 5 000 iterations used a burn-in and the subsequent 10 000 samples thinned every 10th value. This gave 1 000 samples for each replicate chain, 25 000 samples for each ASI chain, and an overall number of 75 000 posterior samples.
1 2233338 13 58377159 25 76049395 37 118468553 49 140867938 61 159791628
2 4137393 14 58382846 26 76077143 38 121618371 50 146419285 62 159804903
3 5038795 15 58405947 27 89904882 39 121620787 51 146579888 63 169461571
4 6104625 16 58512237 28 89982853 40 121632432 52 146594746 64 169492101
5 7255106 17 60095604 29 99905803 41 121646886 53 146601081 65 169497585
6 12531065 18 60098968 30 99906993 42 121664112 54 146601295 66 169508272
7 12746807 19 62645954 31 99914139 43 121715319 55 159501673 67 169512145
8 13052644 20 68324960 32 103387440 44 123131254 56 159582382 68 178995657
9 16224346 21 68328854 33 103388364 45 123925271 57 159728987 69 178997778
10 18998569 22 72187269 34 103394499 46 125880208 58 159732983
11 29378090 23 73444439 35 103404111 47 129083281 59 159767453
12 46880128 24 74270459 36 103433922 48 129084581 60 159780373
Table 4: GWAS example: the genomic distance of each variable chosen using the PIP screening
Appendix D Concrete data
The original variables in the Concrete data are shown in Table 5. The results of using the Agglomerative Method with Prior I and Prior II are shown in Figures 8 and 9 respectively. The credible sets are similar to the corresponding ones using the Thresholding Method but are slightly more conservative.
Original variables $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $X_{5}$ $X_{6}$ $X_{7}$ $X_{8}$ Cement Blast Furnace Fly Water Superplasticizer Coarse Fine Age Slag Ash Aggregate Aggregate
Table 5: Concrete examples original variables
Figure 8: Concrete data: Agglomerative Method with prior I, $M=2$. Black lines indicate the blocks of the partition.
Figure 9: Concrete data: Agglomerative Method with prior II, $M=2$. Black lines indicate the blocks of the partition.
Appendix E ASI algorithm for a linear model with two-way interactions
We are interested in a model with main effects and interactions. Let $X$ be a design matrix for the main effects, $Z$ be a design matrix for the interactions and let $b_{i,1}$ and $b_{i,2}$ be the main effects that correspond to the $i$-th interaction. We assume the linear model
$Y=X_{\gamma}+Z_{\kappa}+\epsilon$
where $\epsilon\sim\mbox{N}(0,\sigma^{2}I)$, $\gamma$ and $\kappa$ are the inclusion variables for the main effects and interactions respectively, $X_{\gamma}$ and $Z_{\kappa}$ are the corresponding design matrices for the included main effects and included interactions respectively. The prior inclusion probability are given a weak heredity prior which has the form
$p(\gamma_{j})=\pi_{M}$
$p(\kappa_{i}=1|\gamma_{b_{i,1}},\gamma_{b_{i,2}})=\left\{\begin{array}[]{ll}\pi_{1}&\mbox{if }\gamma_{b_{i,1}}=0,\gamma_{b_{i,2}}=0\\ \pi_{2}&\mbox{if }\gamma_{b_{i,1}}=0,\gamma_{b_{i,2}}=1\mbox{ or }\gamma_{b_{i,1}}=1,\gamma_{b_{i,2}}=0\\ \pi_{3}&\mbox{if }\gamma_{b_{i,1}}=1,\gamma_{b_{i,2}}=1\end{array}\right.$
where $\pi_{1}<\pi_{2}<\pi_{3}$. Strong heredity (i.e. the idea that an interaction can only be included in the variable if both its main effects are included) corresponds to $\pi_{1}=\pi_{2}=0$. The ASI sampler(Griffin et al., 2021) increasingly struggles to mix as $\pi_{1}$ and $\pi_{2}$ become closer to zero since the posterior distribution will show increasing correlation between interactions and the corresponding main effects.
The ASI sampler can be adapted by introducing a two-step proposal which has the form
$q_{main}(\gamma,\gamma^{\prime})=\prod_{i=1}^{p}A_{i}^{(\gamma_{i}-1)\gamma_{i}^{\prime}}(1-A_{i})^{(\gamma_{i}-1)(\gamma_{i}^{\prime}-1)}D_{i}^{\gamma_{i}(1-\gamma_{i}^{\prime})}(1-D_{i})^{\gamma_{i}\gamma_{i}^{\prime}}$
where $A_{i}$ and $D_{i}$ are the calculated using the Rao-Blackweillised estimator described in Griffin et al. (2021). The interactions are proposed (conditional on $\gamma$ and $\gamma^{\prime}$) from
$q_{int}(\kappa,\kappa^{\prime})=\prod_{i=1}^{q}\tilde{A}_{i}^{(\kappa_{i}-1)\kappa_{i}^{\prime}}(1-\tilde{A}_{i})^{(\kappa_{i}-1)(\kappa_{i}^{\prime}-1)}\tilde{D}_{i}^{\kappa_{i}(1-\kappa_{i}^{\prime})}(1-\tilde{D}_{i})^{\kappa_{i}\kappa_{i}^{\prime}}$
where $\tilde{A}_{i}$ and $\tilde{D}_{i}$ by
$\tilde{A}_{i}=\zeta\min\left\{1,\frac{\tilde{\pi}_{i,\gamma^{\prime}_{b_{i,1}}+\gamma^{\prime}_{b_{i,2}}}}{1-\tilde{\pi}_{i,\gamma_{b_{i,1}}+\gamma_{b_{i,2}}}}\right\}$
and
$\tilde{D}_{i}=\zeta\min\left\{1,\frac{1-\tilde{\pi}_{i,\gamma^{\prime}_{b_{i,1}}+\gamma^{\prime}_{b_{i,2}}}}{\tilde{\pi}_{i,\gamma_{b_{i,1}}+\gamma_{b_{i,2}}}}\right\}.$
The parameter $\zeta$ is common to both proposals. The PIPs for the interactions in the proposal $\tilde{\pi}_{i,j}$ are Rao-Blackwellised estimates of the PIP $p(\kappa_{i}=1|\gamma_{b_{i,1}}+\gamma_{b_{i,2}}=j)$. This allows the proposal for the interaction terms to adapt to the included/excluded main effects. This becomes increasingly important if $\pi_{1}$ and $\pi_{2}$ are close to zero. In the limit, in the case of strong heredity, $p(\kappa_{i}=1|\gamma_{b_{i,1}}+\gamma_{b_{i,2}}<2)=0$ and so must be different from $p(\kappa_{i}=1|\gamma_{b_{i,1}}+\gamma_{b_{i,2}}=2)$. These probabilities will also be very different if $\pi_{1}$ and $\pi_{2}$ are close to zero. The proposal is accepted using the usual Metropolis-Hastings ratio.